OpenStudy (anonymous):

What is the experimental probability of rolling a 5 on a standard die if you had 25 trials and 1 success? 1% 4% 5% 25%

OpenStudy (acxbox22):

1/25=4/100=4%

OpenStudy (anonymous):

Thank You